what is the probability of rolling a number less than or equal to 6 on a regular six-sided die?
Answer #1

Probability (P) is given as the ratio of number of outcome (N) to total possible outcome (NP). That is

P =N/Np

On a dice, we have (1,2,3,4,5 and 6 faces)

Np = Total number of possible outcome

=6

Faces less than or equal to 6 = (1,2, 3,4,5 and 6 faces)

N = 6

Therbefore, P = N/NP =6/6 =1

Answer = 1